calculate the pH of each solution

82 Calculate the pH of eau а) 1.34 х 10,3 м НСОч (b) 0.0211M NaOH (0) 0.0104 м 4Ыr (d) 7.02 X 10.5 M Ba(OH2
Answer #1

a) 1.34X103 m HCIOL [he+] = 1:34 X163 m ph= -log(1.34x103) pl= 2.87 (6) 0.0211m Nach [Hoo]=0.0211m port= -log [H] = -log(0-02